# Deep Joint CSI Estimation-Feedback-Precoding for MU-MIMO OFDM Systems

### v1

@@ II-B Background Description
In contrast, the joint architecture integrates source and channel coding into a unified module, allowing the network to adapt to changing channel conditions during training via an end-to-end approach. This integration effectively mitigates the “cliff effect” typically observed in separate architectures. The joint architecture has been successfully implemented in several of our previous works for the CSI feedback task[13, 18] and the image transmission task [6], demonstrating its efficacy. Consequently, the feedback mechanism in this paper also adopts this DJSCC framework.
III DJSCC-Based MU Joint Channel Estimation, Feedback, and Precoding Design
In the MU MIMO-OFDM scenario, we propose an end-to-end joint estimation, feedback and precoding network (JEFPNet), as shown in Fig. 2, by considering the three processes of downlink CSI estimation, uplink CSI feedback under the DJSCC architecture, and downlink precoding design.
@@ III DJSCC-Based MU Joint Channel Estimation, Feedback, and Precoding Design
$\begin{split}\mathop{{\rm{maximize}}}\limits_{\theta,\psi,\mathscr{P}}&R_{\theta,\psi,\mathscr{P}}=\sum\limits_{k=1}^{K}{\sum\limits_{n_{c}=1}^{N_{c}}{\mathrm{log}_{2}\left({1+\frac{\lVert{\mathbf{h}_{d,n_{c}}^{T}(k)\mathbf{v}_{n_{c}}(k)x_{n_{c}}(k)}\rVert^{2}}{\sum\limits_{\begin{subarray}{c}m\neq k\end{subarray}}{\lVert{\mathbf{h}_{d,n_{c}}^{T}(k)\mathbf{v}_{n_{c}}(m)x_{n_{c}}(m)}\rVert^{2}}+\sigma_{d}^{2}}}\right)}}\\ \text{s.t.}&\quad\left\{{\left\{{\mathbf{v}_{n_{c}}(k)}\right\}}_{{n_{c}}=1}^{N_{c}}\right\}_{k=1}^{K}=\mathcal{D}_{\psi}\left({\left\{{\mathcal{F}_{\mathrm{u},k}{\left({\mathcal{E}_{\theta}{\left({\mathcal{F}_{\mathrm{ce},k}{\left({\mathscr{P}}\right)}},\mathscr{P}\right)}}\right)}}\right\}_{k=1}^{K}}\right)\end{split}$ (19)
@@
The received pilot signal vectors are concatenated in the frequency domain to form the received pilot matrix, denoted as $\mathbf{Y}(k)$ for the $k$-th user. Treating the real and imaginary parts as separate channels, this two-channel received pilot matrix is fed into the encoder module for direct compression. The network architecture of the encoder module is shown in Fig. 3. Here, “$\mathrm{Conv}\medspace F|(W_{1},W_{2})|(S_{1},S_{2})$” indicates a convolutional layer with $F$ filters, a window size is $W_{1}\times W_{2}$, and a stride of $S_{1}\times S_{2}$. “$\mathrm{BN}+\mathrm{LeakyReLU}$” represents a batch normalization (BN) layer followed by the “$\mathrm{LeakyReLU}$” activation function. “$\mathrm{Dense}(2Z)$” denotes a FC layer with $2Z$ neurons. In the encoder module, CSI features are first extracted by the convolutional layer, then reshaped into vectors, and finally compressed into $2Z$-dimensional vectors using an FC layer with $2Z$ neurons. The transmit power is fixed through power normalization and OFDM mapping of the feature vectors before transmission.
@@
In this paper, we do not consider multiple access scenarios. We assume that each user feeds back CSI during a single OFDM symbol, with CSI feedback occurring sequentially for each user without inter-user interference during the feedback phase. The semantic feature vectors obtained from each user’s compressed signal are concatenated into a vector of dimension $2KZ$ at the BS through a “$\mathrm{Concatenation}$” layer after passing through individual channels in real number. Following this, an FC layer recovers the CSI dimension, transforming the vector of dimension $2KZ$ to one of dimension $2KMN_{t}$. A “$\mathrm{Reshape}$” layer then converts this vector into a semantic feature tensor with dimensions $M\times N_{t}\times 2K$, where $2K$ represents the number of channels for this tensor. The feature tensor for each channel can thus be treated as a semantic feature matrix with $M$ rows and $N_{t}$ columns. This MU CSI semantic feature tensor is processed by a convolutional network and a transposed convolutional network to extract features and reconstruct dimensions aligned with the dimensional requirements of precoding design. Finally, the resulting feature tensor, combined with downlink noise power, is fed into the precoding design module to complete the precoding design. The outputs of the precoding design module are the newly designed precoding matrix set for $K$ users as shown in Eq. (12), each with $N_{t}$ antennas and $N_{c}$ subcarriers.
The entire network is trained in an end-to-end manner. Inputs to the network include the downlink full CSI for the CE phase and downlink data transmission phase, the uplink CSI for the feedback phase, and the SNR for each phase. The objective of network optimization is to maximize the downlink spectral efficiency, which can be expressed as Eq. (19)
@@ III-A Feature Extraction and Dimensional Reconstruction
By directly extracting the channel CSI for compression at the UE, the source compression rate is indirectly reduced. At the same time, both parameter storage and computational overhead are decreased, because the source data for compression is changed from full CSI to a received pilot signal with lower dimensionality. Since the compressed pilot signals have contained the required downlink CSI semantic information of users, a feature extraction network at the BS is used to capture MU CSI semantic features, and a dimensional reconstruction network is employed to reconstruct the feature dimensions.
The structure of the feature extraction and dimensional reconstruction networks is shown in Fig. 4. First, for the feature tensor of dimension, $M\times N_{t}\times 2K$, features are extracted using multiple residual blocks composed of convolutional layers, BN layers, and activation function layers. The dimensional reconstruction network employs $I$ transposed convolutional layers to reconstruct low-dimensional features to the target dimensionality of the precoding vector by adjusting the stride $S_{1}$, followed by a convolutional layer to set the number of output feature channels. Overall, the dimensional transformation $M\times N_{t}\times 2K\rightarrow N_{c}\times N_{t}\times 2K$ is achieved by the dimensional reconstruction network.
@@ III-B Precoding Design
In the precoding design module, we account for the correlation between different subcarriers and employ a consistent network architecture and parameters for the precoding design across subcarriers to reduce both the number of training parameters and storage overhead. The input feature tensor is first concatenated along the user and antenna dimensions, then concatenated with the downlink noise power as known information to assist with power allocation. The concatenated vectors for each subcarrier contain the CSI of MU along with the downlink noise power, so precoding vectors for $K$ users are designed using two FC layers and the tanh activation function, as shown in Fig 5. Note that a power normalization process must follow the last activation function layer to meet the power constraints in Eq. (13). Finally, the newly designed precoding matrix is reshaped to form a precoding tensor with dimensions: width corresponding to the number of BS antennas $N_{t}$, length to the number of subcarriers $N_{c}$, and channels to twice the number of users $K$ (representing real and imaginary components). The set of precoding matrices for $K$ users is then obtained by separating this tensor along the channel dimension, which is expressed as $\mathscr{V}$.
Figure 5: Architecture of MU precoding design module. To facilitate understanding, we represent the feature tensor as a set of $N_{c}\times N_{t}\times 2K$ squares. Similarly, the precoding tensor is depicted as $N_{c}\times N_{t}\times 2K$ squares, with different colors used to distinguish the precoding tensor for different users, each characterized by $N_{c}\times N_{t}\times 2$.
@@ IV Experimental results
In this paper, we propose an end-to-end JEFPNet that incorporates three phases of CSI feedback, beginning at the UE and returning to the UE: downlink CSI estimation, uplink CSI feedback, and downlink precoding design. In this section, we demonstrate the effectiveness of the proposed network by comparing our algorithm with traditional algorithms through simulation experiments. We also verify the performance advantages of the end-to-end multi-module joint designed architecture by comparing the joint network architecture with the corresponding separate network architecture. Finally, we evaluate the scalability of the network with respect to varying feedback overhead and the number of supported users.
@@
This paper considers a MU urban macro-cell (UMa) channel scenario to simulate the proposed network architecture. The users work in an FDD system with an uplink frequency of 2.1 GHz and a downlink frequency of 1.9 GHz. An OFDM system with $N_{c}=96$ subcarriers and a 10 MHz bandwidth is employed for both uplink and downlink. The BS is equipped with $N_{t}=32$ logical antenna ports, implemented using $8\times 8$ dual-polarized physical surface array antennas, with the mapping between logical ports and physical antennas realized through a sparse matrix with power constraints. Each UE is equipped with a single omnidirectional antenna. Other antenna settings, user locations, and scenario configurations follow the 3rd Generation Partnership Project (3GPP) TR 38.901 standard [23].
@@
In the downlink CE phase, sparse pilots with equal spacing are utilized, with a pilot interval of $g=4$, corresponding $M=24$. User locations are randomly distributed following the standard, and Quadriga [24] is used to generate the corresponding uplink and downlink CSI for each user. The uplink and downlink CSI of $K$ users are aggregated to form one sample. The dataset comprises 100,000 training samples, 30,000 validation samples, and 20,000 test samples. Training is conducted over 500 epochs, utilizing the Adam optimizer with a batch size of 128. The initial learning rate is set to 0.001 and is reduced by half if the validation loss does not improve for 20 consecutive epochs. The downlink spectral efficiency is calculated with Eq. (16) under a downlink power constraint of $P=1$.
@@ IV-A Separate Architecture
Figure 6: The system model includes pilot design, CSI estimation, CSI feedback, and precoding design within a separate architecture framework. Additionally, network architecture diagrams for some modules not featured in the joint architecture are provided. For different users, the parameters are shared across modules because they operate in the same scatterer environment.
@@
To evaluate the performance differences between the joint and separate architectures, we designed a separate architecture comprising downlink CE, CSI compression, CSI reconstruction, and precoding design tasks, as shown in Fig. 6. At the UE, instead of the joint CE and CSI compression used in JEFPNet, we first employ a CE network to estimate the downlink CSI. Next, a feature extraction network is applied to extract the downlink CSI features in the frequency domain, followed by CSI compression using the same architecture depicted in Fig. 3. The compressed CSI is then transmitted. The details of the CE network and the processing of downlink CSI feature extraction are both illustrated in Fig. 6. For the CE network, we begin by using the learnable parameter matrix $\mathbf{O}_{\mathcal{M}_{m}}\in\mathbb{C}^{L\times N_{t}}$ with the bias vector $\mathbf{b}_{\mathcal{M}_{m}}\in\mathbb{C}^{N_{t}}$ for each subcarrier posted pilots, i.e., $\mathcal{M}_{m}\in\mathcal{M}$ to simulate the linear minimum mean square error (LMMSE) CE, providing the initial estimation result [25]. Similar to the operation of the learnable pilot matrix, the real and imaginary parts of the received pilot signal $\mathbf{Y}(k)$ of the $K$-th user are processed separately, i.e., the real part of the initial estimated downlink CSI of the $\mathcal{M}_{m}$-th subcarrier can be expressed as $\mathcal{R}\big{(}{\mathbf{y}_{\mathcal{M}_{m}}(k)}\big{)}\mathcal{R}\big{(}{\mathbf{O}_{\mathcal{M}_{m}}}\big{)}-\mathcal{I}\big{(}{\mathbf{y}_{\mathcal{M}_{m}}(k)}\big{)}\mathcal{I}\big{(}{\mathbf{O}_{\mathcal{M}_{m}}}\big{)}+\mathcal{R}\big{(}{\mathbf{b}_{\mathcal{M}_{m}}}\big{)}$, and the imaginary part can be expressed as $\mathcal{R}\big{(}{\mathbf{y}_{\mathcal{M}_{m}}(k)}\big{)}\mathcal{I}\big{(}{\mathbf{O}_{\mathcal{M}_{m}}}\big{)}+\mathcal{I}\big{(}{\mathbf{y}_{\mathcal{M}_{m}}(k)}\big{)}\mathcal{R}\big{(}{\mathbf{O}_{\mathcal{M}_{m}}}\big{)}+\mathcal{I}\big{(}{\mathbf{b}_{\mathcal{M}_{m}}}\big{)}$. A transpose convolution layer is then employed to perform upsampling in the frequency domain, followed by several convolution layers and ReLU activation layers for feature extraction and channel reconstruction. As illustrated in Fig. 6, the feature extraction process consists of convolution and transpose convolution layers, along with batch normalization layers and activation functions.
@@
At the BS, the downlink CSI is first reconstructed using a CSI reconstruction network, which employs an architecture similar to the decoder in [13]. The CSI of different users is processed using CSI reconstruction networks with the same parameters. Following CSI reconstruction, precoding vectors are designed based on the reconstructed CSI of the MU.
@@
It is important to note that in the separate architecture, the modules performing distinct tasks are individually trained with optimization objectives tailored to their respective tasks and then integrated for end-to-end performance testing. For the downlink CSI estimation task, a pilot design and CE network is utilized, and its loss function aims to maximize estimation accuracy, represented by

### v2

@@ II-B Background Description
In contrast, the joint architecture integrates source and channel coding into a unified module, allowing the network to adapt to changing channel conditions during training via an end-to-end approach. This integration effectively mitigates the “cliff effect” typically observed in separate architectures. The joint architecture has been successfully implemented in several of our previous works for the CSI feedback task[ADJSCC, JFPNet] and the image transmission task [AI_Image_Process], demonstrating its efficacy. Consequently, the feedback mechanism in this paper also adopts this DJSCC framework.
III DJSCC-Based MU Joint CE, Feedback, and Precoding Design
In the MU MIMO-OFDM scenario, we propose an end-to-end joint estimation, feedback and precoding network (JEFPNet), as shown in Fig. 2, by considering the three processes of downlink CE, uplink CSI feedback under the DJSCC architecture, and downlink precoding design. To ensure that the proposed network adapts to varying numbers of users, a random user activation mask matrix is introduced in the precoding stage.
@@ III DJSCC-Based MU Joint CE, Feedback, and Precoding Design
$\begin{split}\mathop{{\rm{maximize}}}\limits_{\theta,\psi,\mathscr{P}}&R_{\theta,\psi,\mathscr{P}}=\sum\limits_{k=1}^{K_{\max}}{\sum\limits_{n_{c}=1}^{N_{c}}{{\Upsilon_{k}}\mathrm{log}_{2}\left({1+\frac{\lVert{\mathbf{h}_{d,n_{c}}^{T}(k)\mathbf{v}_{n_{c}}(k)x_{n_{c}}(k)}\rVert^{2}}{\sum\limits_{\begin{subarray}{c}m\neq k\end{subarray}}{\lVert{\mathbf{h}_{d,n_{c}}^{T}(k)\mathbf{v}_{n_{c}}(m)x_{n_{c}}(m)}\rVert^{2}}+\sigma_{d}^{2}}}\right)}},\\ \text{s.t.}&\quad\left\{{\left\{{\mathbf{v}_{n_{c}}(k)}\right\}}_{{n_{c}}=1}^{N_{c}}\right\}_{k=1}^{K}=\mathcal{D}_{\psi}\left({\left\{{\mathcal{F}_{\mathrm{u},k}{\left({\mathcal{E}_{\theta}{\left({\mathcal{F}_{\mathrm{ce},k}{\left({\mathscr{P}}\right)}},\mathscr{P}\right)}}\right)}}\right\}_{k=1}^{K}}\right),\end{split}$ (19)
@@
The received pilot signal vectors are concatenated in the frequency domain to form the received pilot matrix, denoted as $\mathbf{Y}(k)$ for the $k$-th user. The real and imaginary parts of the received signal are treated as separate channels, and the received pilot signal—representing the task-oriented feature information and containing noise—is used as the source input to the DJSCC encoder for direct compression. The network architecture of the encoder module is shown in Fig. 3. Here, $\mathrm{Conv}\medspace F|(W_{1},W_{2})|(S_{1},S_{2})$” indicates a convolutional layer with $F$ filters, a window size is $W_{1}\times W_{2}$, and a stride of $S_{1}\times S_{2}$. $\mathrm{BN}+\mathrm{LeakyReLU}$” represents a batch normalization (BN) layer followed by the $\mathrm{LeakyReLU}$” activation function. $\mathrm{Dense}(2Z)$” denotes a FC layer with $2Z$ neurons. The detailed network parameters are specified in the Fig. 3, and the number of layers and the convolutional kernel size are consistent with those commonly adopted in existing CSI feedback studies [CsiNet_plus, ADJSCC, JFPNet]. In the encoder module, CSI features are first extracted by the convolutional layer, then reshaped into vectors, and finally compressed into $2Z$-dimensional vectors using an FC layer with $2Z$ neurons, which corresponds to the feedback bandwidth $Z$ after real-complex transformation. The transmit power is fixed through power normalization and OFDM mapping of the feature vectors before transmission.
@@
In this paper, we do not consider multiple access scenarios. We assume that each user feeds back CSI during a single OFDM symbol, with CSI feedback occurring sequentially for each user without inter-user interference during the feedback phase. The received compressed signals for each user are first dimensionally recovered using a FC layer with $2MN_{t}$ neurons. The downlink CSI semantic feature tensor, with dimensions $M\times N_{t}\times 2$, is then reconstructed through a $\mathrm{Reshape}$” layer. Therefore, the CSI for each user can initially be regarded as a semantic feature matrix of size $M\times N_{t}\times 2$ at the BS. The neuron parameters are configured with reference [ADJSCC] to align with the separation architecture, thereby ensuring a fair comparison. This MU CSI semantic feature tensor is processed by a convolutional network and a transposed convolutional network to extract features and reconstruct dimensions aligned with the dimensional requirements of precoding design. Finally, the resulting feature tensor, combined with downlink noise power and mask matrix, is fed into the precoding design module to complete the precoding design. The outputs of the precoding design module are the newly designed precoding matrix set for $K$ users as shown in Eq. (12), each with $N_{t}$ antennas and $N_{c}$ subcarriers.
The entire network is trained in an end-to-end manner. Inputs to the network include the downlink full CSI for the CE phase and downlink data transmission phase, the uplink CSI for the feedback phase, the SNR for each phase and the mask matrix for determining the activated users. The objective of network optimization is to maximize the downlink spectral efficiency, which can be expressed as Eq. (19). Here, $\Upsilon_{k}$ denotes the $k$-th element of the mask matrix $\mathbf{\Upsilon}\in\mathbb{R}^{K_{\max}}$. Specifically, $\Upsilon_{k}=1$ indicates that the corresponding user is activated, while $\Upsilon_{k}=0$ indicates that the user is inactive. The total number of active users is given by $K=\lVert\mathbf{\Upsilon}\rVert_{1}\leq K_{\max}$.
@@ III-A Feature Extraction and Dimensional Reconstruction
By directly extracting the channel CSI for compression at the UE, both parameter storage and computational overhead are decreased, because the source data for compression is changed from full CSI to a received pilot signal with lower dimensionality. Since the compressed pilot signals have contained the required downlink CSI semantic information of users, a feature extraction network at the BS is used to capture MU CSI semantic features, and a dimensional reconstruction network is employed to reconstruct the feature dimensions.
The structure of the feature extraction and dimensional reconstruction networks is shown in Fig. 4. First, for the feature tensor of dimension, $M\times N_{t}\times 2$, features are extracted using multiple residual blocks composed of convolutional layers, BN layers, and activation function layers. The dimensional reconstruction network employs $I$ transposed convolutional layers to reconstruct low-dimensional features to the target dimensionality of the precoding vector by adjusting the stride $S_{1}$, followed by a convolutional layer to set the number of output feature channels. Overall, the dimensional transformation $M\times N_{t}\times 2\rightarrow N_{c}\times N_{t}\times 2$ is achieved by the dimensional reconstruction network. The size of $I$ is determined by $M$, $N_{c}$, and the upsampling step $S_{1}$, and is typically calculated as $I=N_{c}/(MS_{1})$. If this expression does not yield an integer, $M$ can be adjusted to $M^{\prime}$ to ensure that the reconstructed feature dimensions remain aligned with $N_{c}$.
@@ III-B Precoding Design
In the precoding design module, we account for the correlation between different subcarriers and employ a consistent network architecture and parameters for the precoding design across subcarriers to reduce both the number of training parameters and storage overhead. For the case of a variable number of users, we design a Transformer-based precoding design module utilizing a masked multi-head attention mechanism. The precoding design network for a single subcarrier is also depicted in Fig. LABEL:sub@PrecodingDesign_Transformer. Let $\mathbf{X}\in\mathbb{R}^{K_{\max}\times 2N_{t}}$ denote the features of $K_{\max}$ users on the same subcarrier. The corresponding query, key, and value matrices are expressed as
$\begin{split}&\mathbf{Q}^{(h)}=\mathbf{X}\mathbf{W}_{\mathrm{Q}}^{(h)},\\ &\mathbf{K}^{(h)}=\mathbf{X}\mathbf{W}_{\mathrm{K}}^{(h)},\\ &\mathbf{V}^{(h)}=\mathbf{X}\mathbf{W}_{\mathrm{V}}^{(h)},\end{split}$ (20)
where $\mathbf{W}_{\mathrm{Q}}^{(h)}\in\mathbb{R}^{2N_{t}\times E}$, $\mathbf{W}_{\mathrm{K}}^{(h)}\in\mathbb{R}^{2N_{t}\times E}$, and $\mathbf{W}_{\mathrm{V}}^{(h)}\in\mathbb{R}^{2N_{t}\times E}$ are trainable parameters associated with the $h$-th attention head. $E$ represent embedding dimension of outputs.
After applying the mask, the output of the $h$-th attention layer can be expressed as
$\mathrm{Softmax}\left(\frac{\mathbf{Q}^{(h)}\mathbf{K}^{(h)^{T}}}{\sqrt{K_{\max}}}\odot\left(\mathbf{\Upsilon}\mathbf{\Upsilon}^{T}\right)\right)\mathbf{V}^{(h)},$ (21)
where $\mathrm{Softmax}(\cdot)$ denotes the softmax function, and $\mathbf{\Upsilon}\mathbf{\Upsilon}^{T}$ represents the mask matrix, ensuring that the attention computation is restricted to the activated users.
In our setup, the embedding dimension $E$ is set to 256. Therefore, the embedding of noise power with shape $\mathbb{R}^{256}$ is incorporated as auxiliary information to enhance precoding design. The output of the multi-head attention layer is added with the input features and the noise power embedding, followed by several FC layers to restore each feature dimension to $2N_{t}$, corresponding to the real and imaginary parts of the precoding vector. Finally, the mask matrix selects the precoding vectors of the active users, which are normalized to form the final precoding matrix.
Compared with the DNN-based precoding architecture [JFPNet, YuWei_JEFPNet, MU_Adaptive], which can employs zero-setting operations for user-number adaptation as in [MU_Adaptive] and is illustrated in Fig. LABEL:sub@PrecodingDesign_DNN, our design offers a key advantage: as shown in Eq. (20), the number of trainable parameters remains constant regardless of $K_{\max}$. In [MU_Adaptive], user adaptation is achieved by setting the features of inactive users to zero, which resembles pruning. However, the performance of this approach heavily depends on the depth and width of the following DNN. Specifically, when $K_{\max}$ is large, using a shallow or narrow network results in degraded performance for $K=K_{\max}$, whereas a deep or wide network causes inefficiency and performance loss when $K\ll K_{\max}$. In contrast, the proposed user-adaptive Transformer-based precoding design avoids this trade-off.
When the system is extended to a hybrid beamforming architecture, two precoding design modules should be separately designed for $\mathbf{F}_{\mathrm{RF}}$ and $\mathbf{F}_{\mathrm{BB}}$. It should be noted that the output dimensions of the last layer of the FC for the two networks should be $2N_{\mathrm{RF}}$ and $2N_{t}$, respectively. Furthermore, for analog precoding, the output must be amplitude-constrained to ensure that only the phase is adjusted.
Figure 5: Architecture of MU precoding design module. LABEL:sub@PrecodingDesign_Transformer Transformer precoding method with proposed user-adaptive strategy. LABEL:sub@PrecodingDesign_DNN DNN precoding method [JFPNet] with zeroing user-adaptive strategy [MU_Adaptive]
@@ IV Experimental results
In this paper, we propose an end-to-end JEFPNet with user-number adaptability that integrates three phases of CSI feedback starting at the UE and eventually returning to the UE: downlink CE, uplink CSI feedback, and downlink precoding design. In this section, we demonstrate the effectiveness of the proposed network by comparing our algorithm with traditional algorithms through simulation experiments. We further validate the effectiveness of the DJSCC architecture for joint CE, CSI feedback, and precoding design, as well as the advantages of this joint design over separated designs. In addition, we demonstrate the effectiveness of the proposed user-number-adaptive mechanism through comparative experiments. We also verify the performance advantages of the end-to-end multi-module joint designed architecture by comparing the joint network architecture with the corresponding separate network architecture. Finally, we evaluate the network’s scalability under varying feedback overheads and the impact of the scenario shift problem on its performance. We also examine the network’s compatibility with existing digital communication systems and provide a statistical characterization of the model storage, distribution, and computational overheads.
@@
This paper considers a MU urban macro-cell (UMa) channel scenario to simulate the proposed network architecture. In this paper, we adopt a two-sided model similar to [YuWei_JEFPNet, ADJSCC], where the BS collects channel data during idle periods for offline training and subsequently distributes the trained model. When the UE enters a cell and initiates a communication request, the BS distributes the corresponding encoder model of that cell to the UE. The users work in an FDD system with an uplink frequency of 2.1 GHz and a downlink frequency of 1.9 GHz. An OFDM system with $N_{c}=96$ subcarriers and a 10 MHz bandwidth is employed for both uplink and downlink. The BS is equipped with $N_{t}=32$ logical antenna ports, implemented using $8\times 8$ dual-polarized physical surface array antennas, with the mapping between logical ports and physical antennas realized through a sparse matrix with power constraints. Each UE is equipped with a single omnidirectional antenna. Other antenna settings, user locations, and scenario configurations follow the 3rd Generation Partnership Project (3GPP) TR 38.901 standard [3GPP38901]. For example, for the UMa scenario, the cell radius is set to 250 m, the BS height to 25 m, and the minimum UE-to-BS distance to 35 m. The indoor user ratio is set to 80%. The channel between the $n_{r}$-th receiving antenna and the $n_{t}$-th transmitting antenna can be expressed as
$\begin{split}{H_{{n_{r}},{n_{t}}}}(\tau)&=\sqrt{\frac{1}{{K_{R}+1}}}H_{{n_{r}},{n_{t}}}^{{\rm{NLOS}}}(\tau)\\ &+\sqrt{\frac{K_{R}}{{K_{R}+1}}}H_{{n_{r}},{n_{t}}}^{{\rm{LOS}}}(\tau)\delta(\tau-{\tau_{1}})\end{split}$ (22)
where $\delta(\cdot)$ denotes the Dirac delta function, $K_{R}$ is the Rician $K$-factor determined by the propagation scenario, $T$ represents the cluster delays, $H_{{n_{r}},{n_{t}}}^{{\rm{LOS}}}$ is the line-of-sight (LOS) channel component, and $H_{{n_{r}},{n_{t}}}^{{\rm{NLOS}}}$ is the sum of the non-line-of-sight (NLOS) multipath components.
@@
In the downlink CE phase, sparse pilots with equal spacing are utilized, with a pilot interval of $g=4$, corresponding to $M=24$. User locations are randomly distributed following the standard, and QuaDriGa [QuaDriGa, QuaDriGa_3D] is used to generate the corresponding uplink and downlink CSI for each user. For the UMa scenario, the QuaDriGa configuration is set to “3GPP_38.901_UMa”, with the default parameters applied, including 17 channel clusters. The uplink and downlink CSI of $K$ users are aggregated to form one sample. The dataset comprises 100,000 training samples, 30,000 validation samples, and 10,000 test samples. In practical deployments, datasets can be acquired through idle-time measurements, and the network can be trained offline. Training is conducted over 500 epochs, utilizing the Adam optimizer with a batch size of 128. The initial learning rate is set to 0.001 and is reduced by half if the validation loss does not improve for 20 consecutive epochs. The downlink spectral efficiency is calculated with Eq. (16) under a downlink power constraint of $P=1$.
@@ IV-A Separate Architecture
Figure 6: The system model includes pilot design, CE, CSI feedback, and precoding design within a separate architecture framework. Additionally, network architecture diagrams for some modules not featured in the joint architecture are provided. For different users, the parameters are shared across modules because they operate in the same scatterer environment.
@@
To evaluate the performance differences between the joint and separate architectures, we designed a separate architecture comprising downlink CE, CSI compression, CSI reconstruction, and precoding design tasks, as shown in Fig. 6. At the UE, instead of the joint CE and CSI compression used in JEFPNet, we first employ a CE network to estimate the downlink CSI. Next, a feature extraction network is applied to extract the downlink CSI features in the frequency domain, followed by CSI compression using the same architecture depicted in Fig. 3. The compressed CSI is then transmitted. The details of the CE network and the processing of downlink CSI feature extraction are both illustrated in Fig. 6. The parameter settings of the CE module follow those reported in existing CE studies [CE_Methods], while the feature extraction network is adapted from the analysis transformation network used in CSI feedback research [ADJSCC]. For the CE network, we begin by using the learnable parameter matrix $\mathbf{O}_{\mathcal{M}_{m}}\in\mathbb{C}^{L\times N_{t}}$ with the bias vector $\mathbf{b}_{\mathcal{M}_{m}}\in\mathbb{C}^{N_{t}}$ for each subcarrier posted pilots, i.e., $\mathcal{M}_{m}\in\mathcal{M}$ to simulate the linear minimum mean square error (LMMSE) CE, providing the initial estimation result [CE_Methods]. Similar to the operation of the learnable pilot matrix, the real and imaginary parts of the received pilot signal $\mathbf{Y}(k)$ of the $K$-th user are processed separately, i.e., the real part of the initial estimated downlink CSI of the $\mathcal{M}_{m}$-th subcarrier can be expressed as $\mathcal{R}\big({\mathbf{y}_{\mathcal{M}_{m}}(k)}\big)\mathcal{R}\big({\mathbf{O}_{\mathcal{M}_{m}}}\big)-\mathcal{I}\big({\mathbf{y}_{\mathcal{M}_{m}}(k)}\big)\mathcal{I}\big({\mathbf{O}_{\mathcal{M}_{m}}}\big)+\mathcal{R}\big({\mathbf{b}_{\mathcal{M}_{m}}}\big)$, and the imaginary part can be expressed as $\mathcal{R}\big({\mathbf{y}_{\mathcal{M}_{m}}(k)}\big)\mathcal{I}\big({\mathbf{O}_{\mathcal{M}_{m}}}\big)+\mathcal{I}\big({\mathbf{y}_{\mathcal{M}_{m}}(k)}\big)\mathcal{R}\big({\mathbf{O}_{\mathcal{M}_{m}}}\big)+\mathcal{I}\big({\mathbf{b}_{\mathcal{M}_{m}}}\big)$. A transpose convolution layer is then employed to perform upsampling in the frequency domain, followed by several convolution layers and ReLU activation layers for feature extraction and channel reconstruction. As illustrated in Fig. 6, the feature extraction process consists of convolution and transpose convolution layers, along with batch normalization layers and activation functions.
@@
At the BS, the downlink CSI is first reconstructed using a CSI reconstruction network, which employs an architecture similar to the decoder in [ADJSCC]. The CSI of different users is processed using CSI reconstruction networks with the same parameters. Following CSI reconstruction, precoding vectors are designed based on the reconstructed CSI of the MU.
@@
It is important to note that in the separate architecture, the modules performing distinct tasks are individually trained with optimization objectives tailored to their respective tasks and then integrated for end-to-end performance testing. For the downlink CE task, a pilot design and CE network is utilized, and its loss function aims to maximize estimation accuracy, represented by